OpenStudy (anonymous):

medal and fan Solve the equation. x – 2/7 = 3

OpenStudy (dmndlife24):

x - 2/7 = 3 (Add 2/7 to both sides) x = 3 + 2/7 (Simplify) x = 23/7

OpenStudy (anonymous):

Thanks

OpenStudy (dmndlife24):

No problem